# Predicting Trust In Autonomous Vehicles: Modeling Young Adult Psychosocial Traits, Risk-Benefit Attitudes, And Driving Factors With Machine Learning

### v1

@@ Abstract.
$c$s: Human-centered computing User modelsccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
Potential benefits of Autonomous Vehicles (AVs) range from reduced traffic, less crashes, and lower environmental impact to improvements in efficiency and risk management (Hewitt et al., 2019). However, widespread adoption of AVs is impeded by a lack of rider trust in AV decisions and the real-world consequences of adoption (Kenesei et al., 2022). AVs are a clear case of a more general lack of public trust in Artificial Intelligence (AI), spanning many domains (Bedué and Fritzsche, 2022).
@@
In the present study, we use survey methods (n = 1457) to measure a broader range of variables than any previous research on this topic, bridging a diverse spectrum of driving and non-driving related research. We include variables ranging from cognitive and psychological traits to cultural values, driving style, and technology attitudes. We also include a wide range of specific risks and benefits in AVs, as these may be important contributors to trust perceptions (Ayoub et al., 2021). Including such a wide range of variables allows us to build a more nuanced and comprehensive understanding of the factors that influence trust than has been previously possible. We are able to capture the complex interactions between variables and assess the relative importance of the factors in our analysis by using machine learning to predict trust in AVs, and then applying SHapley Additive eXplanations (SHAP) (Lundberg et al., 2020) to “peek beneath the hood” and assess how important different factors were to our models’ predictions. Utilizing these advanced analysis methods enhances predictability and builds deeper insights into how a wide range of traits relate to trust judgements.
@@
In Summary, we contribute:
@@
The work presented here builds new understanding on how personal factors can be used to predict trust in AVs, a crucial step towards designing future, more trustworthy human-AV interactions.
@@ 3. Method
In the present study, we use survey methods to assess a wide range of personal factors from a pool of young adults. Then, we build models predicting a person’s trust in autonomous vehicles from those factors, using a variety of machine learning techniques. Following, we use SHAP (Lundberg et al., 2020) to make the top performing models explainable, giving us measures of feature importance. We add additional nuance to our study with ablation experiments knocking out specific feature groups from the models in order to derive greater insight into how factors contribute to AV trust. Figure 1 describes the overall modeling process. To look at trust differences between the “high” and “low” groups at the level of each factor, we also perform a more traditional descriptive analysis using Kruskal-Wallis tests as a means to provide comparison to our ML results.
Figure 1. Overall process architecture, moving from data collection and processing to model development and explanations using SHAP.
[…]
3.1. Participants and Data Collection
@@
3.2. Survey Design
3.2.1. Survey Scales and Questionnaires
The survey was designed to be as comprehensive as possible while considering limits on survey duration and participant attention. Table 2 shows factors related to driving-specific attitudes and behaviors, Table 3 shows demographic, psychosocial, and general attitude measures, and Table 4 details risk and benefit measures as well as the trust measures used to form our composite trust outcome. Given the large number of variables, variable descriptions that are not relevant to the reported results are omitted from these tables; all variable names were designed to be descriptive and intuitive.
3.2.2. Trust as an Outcome
@@
Explanation of measures used in the study: Risks, Benefits, and Trust Category Factor Description Perceptions AV Feasibility The AV and the infrastructure necessary to use the AV are practically feasible. Overall Risk-Benefit How do you perceive the risk-benefit trade-off of using an autonomous vehicle?             [Scale: All Risk, No Benefit- No Risk, All Benefit (1-7)] Mental Model Autonomous vehicles make decisions like humans do. Benefits Reduce Accidents Autonomous vehicles can reduce traffic crashes. Reduce Traffic Autonomous vehicles can reduce traffic congestion. Reduce Emissions Autonomous vehicles can reduce vehicle emissions and pollution. Improve Fuel Economy Autonomous vehicles can improve fuel economy. Reduce Transport Cost Autonomous vehicles can reduce transport costs. Improve Mobility Autonomous vehicles can increase mobility for those who are unable to drive. Free Time Autonomous vehicles will give me more free time (e.g. to text, play games). Increase Fun An autonomous vehicle would make driving more interesting or fun. Improve Efficiency Using an autonomous vehicle would enable me to reach my destination quickly. Risks System Failure I am concerned about equipment and system failures in autonomous vehicles. Legal Liability I am concerned about the legal liability of drivers or owners of autonomous vehicles. Hacking I am concerned that the computer systems of autonomous vehicles may be hacked. Data Privacy I am concerned about sharing driving data with a company or government agency. Performance (varied) I am concerned that autonomous vehicles cannot handle varied weather and terrain. Cost I am concerned that autonomous vehicles will be too expensive. Losing Control I am concerned about giving up driving control to an autonomous vehicle. Lacking Understanding I am afraid that I would not understand an autonomous vehicle’s decisions. Ease of Use I would find an autonomous vehicle easy to use. Trust Dependability Autonomous vehicles are more dependable than human drivers. (Outcome) Adaptability Autonomous vehicles will adapt well to new environments. Goal Alignment The goals for autonomous vehicles are aligned with my own. Danger to Self I believe that using an autonomous vehicle would be dangerous to me. Danger to Others I believe that autonomous vehicles would be dangerous to other drivers/pedestrians. Safety I would feel safe while using an autonomous vehicle. General Trust I would trust an autonomous vehicle. Good Idea Using an autonomous vehicle would be a good idea. Positivity Overall, I feel positive about autonomous vehicles. Recommendation I will recommend family members and friends to ride in autonomous vehicles. All questions across categories were adapted from (Hewitt et al., 2019; Fagnant and Kockelman, 2015; Acheampong and Cugurullo, 2019; Jian et al., 2000; Zhang et al., 2023). Unless otherwise specified, scale: Strg. Disagree- Strg. Agree (1-5).
@@ 3.3. Model Development
After deriving an exploratory understanding of the dataset and its features using descriptive statistics and Kruskal-Wallis tests (discussed in Section 4.1), we investigated the potential for machine learning systems to predict users trust in AVs. We discuss our approach to designing such a model below.
@@
Since this data comes from surveys, we took multiple steps to achieve a concise feature set and mitigate the possibility that some features could be correlated with each other. Following recent work around modeling personality traits (Saha et al., 2021), we first filtered features using their Variance Inflation Factor (VIF) with respect to the raw trust scores, using a threshold of 10. This helped us detect and reduce multicollinearity. To further reduce multicollinearity, next we ran feature selection through Lasso regression (Tibshirani, 1996) on the raw trust scores. This resulted in a final set of 130 features, which was used for all our machine learning experiments and is henceforth referred to as the full feature set.
@@
Next, we applied several classical machine learning models on the derived feature set in order to find the best performing method. Given our task of binary classification (“low” vs. “high” trust), we used models similar to (Bedmutha et al., 2024) to achieve a balance of interpretability and performance. Specifically, we applied Logistic Regression (LR), Support Vector Machines (SVM) with linear and radial kernels, Decision Tree Classifiers (DTC), Random Forest (RF) and Gradient-Boosted Decision Trees. We used XGBoost implementation (Chen and Guestrin, 2016) for the gradient-boosted trees.
@@ 4.3. Ablation Studies: Systematic Isolation and Elimination of Risk-Benefit Factors
Table 7. Summary of the three experiments on our dataset. The full feature set showed the best performance, followed closely by the feature subset (only risk and benefit features). The full feature set without risk and benefit features performed substantially worse.
[…]
@@ 4.3.2. Full Model Without Risk-Benefit Features
Though 73.8% is not excellent predictive accuracy, from a theoretical lens it is still predictive enough to draw initial insights on how non-risk-benefit features impact a person’s judgement of high and low trust. We found that Trust in Tech Companies, Trust in Automakers, AV Feasibility, Know AV Purpose, and Trust in Government Authorities were the strongest predictors (Fig. 6); higher values of each were associated with higher trust (see Fig. 7).
[…]
Figure 7. Full Feature Model without risk/benefit factors: SHAP summary plot showing feature value impact on trust. This plot shows how the value of a feature impacts the model’s output. The more extreme the SHAP value, the more indicative that value was of being high trust (positive SHAP values) or low trust (negative SHAP values). Features are organized by overall importance.
[…]
@@ 5. Discussion
A general principle that resulted from our study is that perceptions of risks and benefits are the most important factors for predicting trust. In our full feature model, 12 of the 20 most important features (based on mean SHAP value) were related to a person’s assessment of the risks and benefits of using AVs. Using ablation studies, the importance of risk-benefit perceptions was confirmed: risk and benefit factors alone had similar predictive ability as our full model, whereas a knockout model without risk and benefit factors performed far worse. The importance of risks and benefits has been highlighted in the past (Ayoub et al., 2021; Hewitt et al., 2019; Zhang et al., 2023), however, the present study is the first to assess the relative importance of specific risk and benefit perceptions compared to other types of personal factors for predicting trust.
@@
Psychosocial and many driving and technology-related factors were far less important than expected — In Section 2.1, we discussed a plethora of prior work suggesting that psychosocial factors – including personality, self-esteem, risk perception, and need for control – will impact a person’s trust judgement for AVs. This did not appear to be the case for our model’s prediction of trust. Further, exempting the few factors noted above, technology and driving factors – including driving style, driving cognitions, and technology self-efficacy – also showed little impact.
@@ 5.1. Implications for Trustworthy Human-AV Interaction Design and Research
Explainable AI and Transparency — The high importance of concerns around ease of use, giving up control, and lack of understanding around AV decision-making highlights the critical need for continued research and design efforts on AV explainability and explainable AI more generally. From a design perspective, our results echo prior work on the importance of XAI (see Section 2.1), particularly for research that supports clear, real-time explanations for AV decisions. A special focus should be put on XAI explanations that are understandable by riders from diverse communities with a wide range of backgrounds and expertise levels. Failure to satisfy concerns over ease of use, vehicle performance, and safety and privacy may result in refusal to ride with an AV.
@@
Policy and Regulation Implications — Given the importance of institutional trust (in technology companies, automakers, and government agencies), our results suggest that clear and consistent regulatory standards are crucial for building public trust in AVs. Policymakers should focus on establishing transparent safety protocols, rigorous testing standards, certification processes, and data sharing policies that are easy for the public to understand. Clear regulations can reassure the public that AVs meet high safety standards and are regularly audited for compliance. These are similar to the recommendations made by Fagnant and Kockelman (2015), though current realization of these recommendations leaves much to be desired in this regard. Our work highlights transparency in regulation as a way to increase trust, particularly for cases when a person has low trust in institutions in general. Broader regulation focused on the most important areas of risk discussed previously – vehicle performance, usability, and security – should be prioritized.

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713188isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing User modelsccs: Human-centered computing HCI theory, concepts and models
@@ 1. Introduction
Potential benefits of Autonomous Vehicles (AVs) range from reduced traffic, less crashes, and lower environmental impact to improvements in efficiency and risk management (Hewitt et al., 2019). However, widespread adoption of AVs is impeded by a lack of rider trust in AV decisions and the real-world consequences of adoption (Kenesei et al., 2022). AVs are a clear case of a more general lack of public trust in Artificial Intelligence (AI), spanning many domains (Bedué and Fritzsche, 2022). Distrust in AI is particularly evident for opaque “black box” systems, whose decision-making processes are not human accessible (Gunning and Aha, 2019; Miller, 2019).
It has long been established that trust is a major determinant of adoption of AI-based technologies like AVs, where people’s beliefs about a system directly inform their willingness to accept and rely on the system (Lee and See, 2004; Muir, 1994). The relationship between trust and adoption may be especially strong in high-stakes contexts like driving, where users must rely on the technology for personal safety (Hoff and Bashir, 2015). Highly influential models by Choi and Ji (2015) and Hoff and Bashir (2015) demonstrate how trust is a necessary pre-condition for intention to use autonomous vehicles: if a person does not trust an AV, they will not be willing to ride in it. Liao and Sundar (2022) discuss how trust impacts users’ perception of risk and shapes their interactions with AI-based systems, meaning that even a perfectly safe AV may face adoption challenges if users do not trust it. In essence, this means that no matter how safe or high performing an AV is, without trust, adoption efforts will fail.
Trust reflects the subjective, psychological confidence that users place in the technology, and does not automatically follow when a system is high-performing or safe (Muir, 1994). In this way, trust serves as a bridge between vehicle performance and user acceptance – without this bridge, AVs may not reach widespread use. Understanding user trust perceptions in AV research (as opposed to solely improving AV safety or performance) enables designers and researchers to address specific user perceptions and concerns that may hinder adoption, including those that move beyond driving performance (Kraus et al., 2021). By identifying which personal factors most influence trust, this study offers insights on the foundation of AV trust perceptions, informing the design of AVs that align with user needs beyond technical safety alone.
@@
In the present study, we use survey methods (n = 1457) to measure a broader range of variables than any previous research on this topic, bridging a diverse spectrum of driving and non-driving related research. We include variables ranging from cognitive and psychological traits to cultural values, driving style, and technology attitudes. We also include a wide range of specific risks and benefits in AVs, as these may be important contributors to trust perceptions (Ayoub et al., 2021). Including such a wide range of variables allows us to build a more nuanced and comprehensive understanding of the factors that influence trust than has been previously possible.
We are able to capture the complex interactions between variables and assess the relative importance of the factors in our analysis by using machine learning (ML) to predict trust in AVs. The rationale for using ML as the basis of our inquiry stems from its ability to capture complex, nonlinear interactions among our chosen personal traits, attitudes, and trust that traditional regression methods may miss (Wang et al., 2020). Yarkoni and Westfall (2017) argue that, unlike linear models which assume independent, additive effects, ML can reveal nuanced, multidimensional relationships and should be opted for in research where predictors have the potential to be interrelated, such as in the work presented here. ML models also handle high-dimensional data very effectively, addressing multicollinearity and retaining diverse features better than traditional regression, without overfitting (Breiman, 2001). By using ML, we can retain a comprehensive set of trust-related features without compromising model validity, offering a more holistic view of a wide range of factors potentially impacting trust. We apply SHapley Additive eXplanations (SHAP) (Lundberg et al., 2020) to “peek beneath the hood” of our models, adding interpretability to the predictive power of our ML models by showing precisely how important different factors were to predictions. This transparency makes our ML insights actionable, guiding AV design and policy toward strategies that directly address users’ trust-building needs. By utilizing these advanced analysis methods, we can build deeper insights into how a wide range of traits relate to trust judgements.
@@
The work presented here builds new understanding on how personal factors can be used to predict trust in AVs, a crucial step towards designing future, more trustworthy human-AV interactions. In Summary, we contribute:
@@ 2.1. Personal Factors Impacting Trust in AVs
Figure 1. Overall process architecture, from data collection and processing to model development and explanations using SHAP.
[…]
@@ 3. Method
In the present study, we use survey methods to assess a wide range of personal factors from a pool of young adults. Then, we build models predicting a person’s trust in autonomous vehicles from those factors, using a variety of machine learning techniques. Following, we use SHAP (Lundberg et al., 2020) to make the top performing models explainable, giving us measures of feature importance. We add additional nuance to our study with ablation experiments knocking out specific feature groups from the models in order to derive greater insight into how factors contribute to AV trust. Figure 1 describes the overall modeling process. Similar to concurrent work on clustering (Bedmutha et al., 2024a; Saha et al., 2021), to look at trust differences between the “high” and “low” groups at the level of each factor, we also perform a more traditional descriptive analysis using Kruskal-Wallis tests as a means to provide comparison to our ML results.
@@
3.1. Participants and Data Collection
@@
Explanation of measures used in the study: Risks, Benefits, and Trust Category Factor Description Perceptions AV Feasibility The AV and the infrastructure necessary to use the AV are practically feasible. Overall Risk-Benefit How do you perceive the risk-benefit trade-off of using an autonomous vehicle?
[Scale: All Risk, No Benefit- No Risk, All Benefit (1-7)]
@@
Mental Model Autonomous vehicles make decisions like humans do. Benefits Reduce Accidents Autonomous vehicles can reduce traffic crashes. Reduce Traffic Autonomous vehicles can reduce traffic congestion. Reduce Emissions Autonomous vehicles can reduce vehicle emissions and pollution. Improve Fuel Economy Autonomous vehicles can improve fuel economy. Reduce Transport Cost Autonomous vehicles can reduce transport costs. Improve Mobility Autonomous vehicles can increase mobility for those who are unable to drive. Free Time Autonomous vehicles will give me more free time (e.g. to text, play games). Increase Fun An autonomous vehicle would make driving more interesting or fun. Improve Efficiency Using an autonomous vehicle would enable me to reach my destination quickly. Risks System Failure I am concerned about equipment and system failures in autonomous vehicles. Legal Liability I am concerned about the legal liability of drivers or owners of autonomous vehicles. Hacking I am concerned that the computer systems of autonomous vehicles may be hacked. Data Privacy I am concerned about sharing driving data with a company or government agency. Performance (varied) I am concerned that autonomous vehicles cannot handle varied weather and terrain. Cost I am concerned that autonomous vehicles will be too expensive. Losing Control I am concerned about giving up driving control to an autonomous vehicle. Lacking Understanding I am afraid that I would not understand an autonomous vehicle’s decisions. Ease of Use I would find an autonomous vehicle easy to use. Trust Dependability Autonomous vehicles are more dependable than human drivers. (Outcome) Adaptability Autonomous vehicles will adapt well to new environments. Goal Alignment The goals for autonomous vehicles are aligned with my own. Danger to Self I believe that using an autonomous vehicle would be dangerous to me. Danger to Others I believe that autonomous vehicles would be dangerous to other drivers/pedestrians. Safety I would feel safe while using an autonomous vehicle. General Trust I would trust an autonomous vehicle. Good Idea Using an autonomous vehicle would be a good idea. Positivity Overall, I feel positive about autonomous vehicles. Recommendation I will recommend family members and friends to ride in autonomous vehicles.
All questions across categories were adapted from (Hewitt et al., 2019; Fagnant and Kockelman, 2015; Acheampong and Cugurullo, 2019; Jian et al., 2000; Zhang et al., 2023). Unless otherwise specified, scale: Strg. Disagree- Strg. Agree (1-5).
3.2. Survey Design and Administration
The survey was designed to be as comprehensive as possible while considering limits on survey duration and participant attention. Our questions were heavily based on validated questionnaires and survey measures to enhance the reliability and interpretability of responses; we relied on previously tested measures to ensure that our questions were well-aligned with established constructs and our own expectations. Given the use of validated questions, we deemed that a formal cognitive pretest was not necessary. However, we conducted a pretest with approximately 10 participants and qualitative feedback on the survey was solicited to ensure that questions were straightforward, clear, free of errors, and not too burdensome. The results from these pilot participants was not included in the final analysis. The pretest allowed us to refine any questions that appeared ambiguous or confusing to participants. The survey itself can be found in Supplementary Material.
Participants were provided with a brief description of autonomous vehicles at the beginning of the survey, specifying AVs as “vehicles that use technology, such as artificial intelligence and special cameras, to drive without the need for human intervention. They sense their environment, adjusting speed and direction to bring a passenger from one location to another safely.” This was accompanied by a simple and clear cartoon visualization of an autonomous sedan in an everyday driving scenario (commuting amongst other vehicles). This introduction aimed to establish the context of interest as AVs focused on typical, day-to-day applications, as opposed to specialized uses like autonomous trucking, racing, or industrial AVs. We chose this commuting-oriented context to ensure that participants were evaluating AVs intended for personal, everyday driving tasks. By avoiding references to specific brands or technical specifications, we allowed participants to respond based on their own knowledge and attitudes toward AVs used in common, relatable scenarios. This approach was designed to capture perceptions of AVs in the specific context of everyday driving, reflecting a more realistic and relevant use case for general public adoption.
3.2.1. Survey Scales and Questionnaires
Table 2 shows factors related to driving-specific attitudes and behaviors, Table 3 shows demographic, psychosocial, and general attitude measures, and Table 4 details risk and benefit measures as well as the trust measures used to form our composite trust outcome. Given the large number of variables, variable descriptions that are not relevant to the reported results are omitted from these tables; all variable names were designed to be descriptive and intuitive.
3.2.2. Trust as an Outcome
@@ 3.3. Model Development
After deriving an exploratory understanding of the dataset and its features using descriptive statistics and Kruskal-Wallis tests (discussed in Section 4.1), we investigated the potential for machine learning systems to predict users trust in AVs, as these advanced modeling methods provide a powerful and nuanced way to derive deep insights from complex data (Wang et al., 2020; Yarkoni and Westfall, 2017). We discuss our approach to designing such a model below.
@@
Since this data comes from surveys, we took multiple steps to achieve a concise feature set and mitigate the possibility that some features could be correlated with each other. Following recent work around modeling personality traits (Saha et al., 2021), we first filtered features using their Variance Inflation Factor (VIF) with respect to the raw trust scores, using a threshold of 10. This helped us detect and reduce multicollinearity. To further reduce multicollinearity, next we ran feature selection through Lasso regression (Tibshirani, 1996) on the raw trust scores. Lasso was opted for because it not only regularizes the model but also performs feature selection by driving less important coefficients to zero, making it more effective than ridge regression for identifying relevant predictors in high-dimensional datasets (Tibshirani, 1996). This resulted in a final set of 130 features, which was used for all our machine learning experiments and is henceforth referred to as the full feature set.
[…]
@@
Next, we applied several classical machine learning models on the derived feature set in order to find the best performing method. Given our task of binary classification (“low” vs. “high” trust), we used models similar to Bedmutha et al. (2024b) and Ayoub et al. (2021) to achieve a balance of interpretability and performance. Specifically, we applied Logistic Regression (LR), Support Vector Machines (SVM) with linear and radial kernels, Decision Tree Classifiers (DTC), Random Forest (RF), Naive Bayes, and Gradient-Boosted Decision Trees. We used XGBoost implementation (Chen and Guestrin, 2016) for the gradient-boosted trees.
@@ 4.2.2. Explaining the Full Feature Model using SHAP
Table 7. Summary of the three experiments on our dataset. The full feature set showed the best performance, followed closely by the feature subset (only risk and benefit features). The full feature set without risks and benefits performed far worse.
[…]
@@ 4.3.1. Feature Subset: Risk-Benefit Features Only
[…]
Figure 7. Full Feature Model without risk/benefit factors: SHAP summary plot showing feature value impact on trust. This plot shows how the value of a feature impacts the model’s output. The more extreme the SHAP value, the more indicative that value was of being high trust (positive SHAP values) or low trust (negative SHAP values). Features are organized by importance.
[…]
@@ 4.3.2. Full Model Without Risk-Benefit Features
Though 73.8% is not excellent predictive accuracy, from a theoretical lens it is still predictive enough to draw initial insights on how non-risk-benefit features impact a person’s judgement of high and low trust. We found that Trust in Tech Companies, Trust in Automakers, AV Feasibility, Know AV Purpose, and Trust in Government Authorities were the strongest predictors (Fig. 6); higher values of each were associated with higher trust (Fig. 7).
@@ 5. Discussion
A general principle that resulted from our study is that perceptions of risks and benefits are the most important factors for predicting trust. The importance of risks and benefit evaluation for trust and acceptance has been highlighted in the past, especially for high-stakes contexts like AVs (Ayoub et al., 2021; Hewitt et al., 2019; Zhang et al., 2023). Unlike prior studies, however, our analysis allowed us to isolate specific types of AV-related risks and benefits, such as individual concerns over system failures and usability. As such, the present study is the first to articulate the relative importance of these risk and benefit perceptions compared to other types of personal factors for predicting trust. In our full feature model, 12 of the 20 most important features (based on mean SHAP value) were related to a person’s assessment of the risks and benefits of using AVs. Using ablation studies, the importance of risk-benefit perceptions was confirmed: risk and benefit factors alone had similar predictive ability as our full model, whereas a knockout model without risk and benefit factors performed far worse.
@@
Psychosocial and many driving and technology-related factors were far less important than expected — In Section 2.1, we discussed a plethora of prior work suggesting that psychosocial factors – including personality, self-esteem, risk perception, and need for control – will impact a person’s trust judgement for AVs. Our results generally contrast with these prior studies, as many of these factors did not appear to be important for our model’s prediction of trust. Further, exempting the few factors noted above, technology and driving factors – including driving style, driving cognitions, and technology self-efficacy – also showed little impact.
@@ 5.1. Implications for Trustworthy Human-AV Interaction Design and Research
Explainable AI and Transparency — The high importance of concerns around ease of use, giving up control, and lack of understanding around AV decision-making highlights the critical need for continued research and design efforts on AV explainability and explainable AI (XAI) more generally. Our results echo prior work on XAI, emphasizing the necessity for explainability contexts where transparency can mitigate user concerns over safety and control (see Section 2.1). For example, these results support the design and study of clear, real-time explanations for AV decisions. A special focus should be put on XAI explanations that are understandable by riders from diverse communities with a wide range of backgrounds and expertise levels. Failure to satisfy concerns over ease of use, vehicle performance, and safety and privacy may result in refusal to ride with an AV.
@@
Implications for Young Adults and Generalizability to Other Populations — Though this study focused on young adults, it is likely that many insights may generalize to broader populations of interest. Certain features – such as ease of use, transparency, and institutional trust – were likely predictive for young adults due to their high exposure to and reliance on technology in daily life. As a generation that has grown up with digital and AI-based tools, prior work has demonstrated that young adults are more accustomed to the presence of autonomous systems and may have greater baseline expectations around usability, control, and information transparency (Anderson and Rainie, 2018). Moreover, young adults may rely on a combination of brand reputation, ethical standards, and transparency practices when evaluating AV trustworthiness (Twenge et al., 2014), potentially explaining why institutional trust emerged as a strong predictor for this group. Focusing on young adults in isolation may also explain the lack of importance of psychosocial, driving, and technology-related factors compared to prior studies on more general populations. Future work segmenting and comparing population clusters could help clarify whether this reduced importance of psychosocial factors is specific to young adults, or if it reflects limitations in the generalizability of prior work over diverse user groups. Though these are potential explanations for why certain findings in this study were found specifically for a young adult population, preference for usability, transparency, as well as many risk and benefit factors are likely important to a wide number of populations. Though future research is necessary to support this claim, we have no reason to suspect that concerns over safety or losing control, for example, would only be relevant to young adults. However, it is probable that different predictors (or at least, different relative importance of predictors) would emerge for other age groups or cultural contexts (Hancock et al., 2011).
Policy, Regulation, and Ethical Implications — Given the importance of institutional trust (in technology companies, automakers, and government agencies), our results suggest that clear and consistent regulatory standards are crucial for building public trust in AVs. Policymakers should focus on establishing transparent safety protocols, rigorous testing standards, certification processes, and data sharing policies that are easy for the public to understand. Clear regulations can reassure the public that AVs meet high safety standards and are regularly audited for compliance. These are similar to the recommendations made by Fagnant and Kockelman (2015), though current realization of these recommendations leaves much to be desired in this regard. Our work highlights transparency in regulation as a way to increase trust, particularly for cases when a person has low trust in institutions in general. Broader regulation focused on the most important areas of risk discussed previously – vehicle performance, usability, and security – should be prioritized.
The development of industry guidelines and regulatory oversight can be a practical way to promote the ethical application and use of trust insights for AV technology. Understanding the individual needs of potential AV users can allow equitable adoption of and access to AV technology. While our research aims to improve user-centered design by identifying factors that predict trust, insights on specific needs could be used for targeted marketing without parallel investments in AV safety and reliability. Such practices could result in users adopting AVs based on trust-building marketing tactics rather than well-informed, safety-centered choices. To prevent misuse, we recommend responsible application of trust-predictive insights, emphasizing transparency, user autonomy, and informed choice. Beyond institutional regulation, companies should ensure that trust-building efforts go hand-in-hand with clear and accessible communication about AV capabilities and limitations. For example, AV manufacturers could provide detailed information on independent safety evaluations, test outcomes, and known limitations of their systems, such as performance in inclement weather or in complex urban settings. This transparency helps users form realistic mental models of AVs, enabling them to trust based on accurate expectations rather than solely on positive marketing.